2 spheres of radii 3cm and 8cm are charged with 4×10(-9) (10 to the power-9) and 16 ×10(-9) (10 to the power -9)C of electricity respectively . If they are joined by a wire how much charge will flow from one to the other.

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

After connecting the two spheres with wire, the potential of both spheres will become equal. Let the charge on the sphere of radiir_1,r_2beq_1,q_2.
Then\dfrac{kq_1}{r_1}=\dfrac{kq_2}{r_2} \implies \dfrac{q_1}{q_2}=\dfrac{8}{3}
Also from conservation of energy,

q_1+q_2=(4\times 10^{-9})+(16\times 10^{-9})=2\times 10^{-8}C
Solving the two equations will give the final charge on each sphere. Subtracting the initial value of charge from it will give the amount of charge flown into or from the other.